# MoA: Heterogeneous Mixture of Adapters for Parameter-Efficient Fine-Tuning of Large Language Models

### v1

@@ Abstract
Jie Cao1, Tianwei Lin 1, Hongyang He, Rolan Yan 2, Wenqiao Zhangthanks: Corresponding author 1, Juncheng Li 1, Dongping Zhang , Siliang Tang 1, Yueting Zhuang 1 1 Zhejiang University, 2 Tencent
@@ 1 Introduction
The rapid development of Large Language Models (LLMs) (Achiam et al., 2023; Yang et al., 2025; Grattafiori et al., 2024a, a) is reshaping the field of NLP, demonstrating cross-domain generalization capabilities across a wide range of tasks. However, as model sizes continue to grow, the computational, storage, and deployment costs of traditional full fine-tuning methods increasingly become bottlenecks in practical applications. As a result, parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; He et al., 2022; Li and Liang, 2021; Lester et al., 2021; Zhang et al., 2024; Han et al., 2024; Lin et al., 2025) are emerging as a key research direction for adapting LLMs. These methods introduce a small number of trainable, lightweight adapter modules on top of frozen pre-trained weights. By doing so, they effectively guide the activation and adjustment of the model’s existing knowledge representations while maintaining low overhead, enabling fast transfer and adaptation to specific downstream tasks.
@@ 3.2 Sparse MoA
As the core objective of PEFT methods is to fully leverage LLMs’ pre-trained knowledge and capabilities via adapters for downstream tasks (Ghosh et al., 2024), Sparse MoA enhances this efficiency by adaptively activating diverse combinations of heterogeneous experts. Moreover, Sparse MoA dynamically adjusts the number of participating experts based on the contextual importance of tokens, thereby mitigating redundant expert computations. For instance, in extreme cases where certain tokens require no expert intervention and can be processed solely by the frozen LLM, Sparse MoA deactivates all experts, thus maximizing computational efficiency.
@@ 4.1 Datasets
We conduct extensive experiments and analyses on both mathematical and commonsense reasoning tasks. For mathematical reasoning, we evaluate performance on six test datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MultiArith Roy and Roth (2016), AddSub Hosseini et al. (2014), AQuA Ling et al. (2017), and SingleEq Koncel-Kedziorski et al. (2015). The training dataset is Math14K constructed by Hu et al. (2023), which includes the training sets of GSM8K and AQuA, augmented with rationales generated by ChatGPT and GPT-4. For commonsense reasoning, we test on eight datasets: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag, WinoGrande Sakaguchi et al. (2021), ARC Challenge (ARC-c), ARC Easy (ARC-e) Chollet (2019), and OBQA Mihaylov et al. (2018). The training dataset for commonsense reasoning, Commonsense15K, is sampled from Commonsense170K Hu et al. (2023) which comprises eight training sets.
@@ 4.2 Baseline Models
To validate the efficacy of our method, we compare it against: (i) the standard LoRA(Hu et al., 2021); (ii) SoTA soft-weighted parameter-efficient MoE methods: MoLoRA Zadouri et al. (2023) and HydraLoRA Tian et al. (2024); and (iii) SoTA sparse parameter-efficient MoE approaches: MoLA (Gao et al., 2024) and AdaMoLE Liu and Luo (2024).
@@ 4.3 Experiment Settings
All experiments use the LLaMA-3.1 8B model (Grattafiori et al., 2024b). To ensure fair comparisons, we apply identical configurations to all LoRA-based methods: LoRA is inserted into four weight matrices in the self-attention module ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$) and one in the FFN module ($W_{\text{up}}$), with rank 8 and alpha 8. For all MoE-LoRA baselines, we set the number of experts to 8. For the original LoRA baseline, we increase the rank to 16 to match the number of trainable parameters across methods. MoA integrates 7 heterogeneous experts, including five LoRA modules, parallel adapters in the FFN layer, and a zero-initialized prompt-tuning module. The bottleneck dimension of the parallel adapter is set to 16, and the prompt length to 10. We use the AdamW optimizer with a learning rate of 6e-3. For commonsense reasoning, input sequences are truncated to 200 tokens; for mathematical reasoning, to 300 tokens. All models are trained with a batch size of 32 on 2×A6000 GPUs.
@@ 5.1 Ablation Study on Soft MoA Components
[…]
Composition Softmax R 84.16 80.21
Composition Sigmoid R 84.96 81.51
@@
The superior performance of LoRA over Prompt-tuning and Parallel Adapter, combined with its inherent inclusion of positionally diverse expert modules, motivated our extension of LoRA’s multi-module experts with a token-level router. While this adaptation yielded significant improvements over standalone LoRA (e.g., +1.32 on mathematical reasoning), it underperformed the full MoA (-1.01), highlighting two critical insights: 1. MoA’s primary strength derives from leveraging multiple LoRA modules as distinct experts. 2. The distinct mechanisms of Prompt Tuning and Parallel Adapter further complement these LoRA-based experts.
5.2 Effect of Threshold Function
  Max Gamma Threshold Function Math
[…]
@@ 6 In-depth Analysis
Figure 3: Comparison of router weight distributions between Soft MoA and AdaMoLE under different random seeds. The MoA method exhibits strong consistency, whereas AdaMoLE does not.
@@ 6.2 Difference between MoA and MoE
As shown in Figure 3, we compare the expert weight distributions across 32 Transformer layers on 50 test samples for Soft MoA and AdaMoLE, trained with different random seeds. Soft MoA exhibits highly consistent expert weight distributions across seeds, as illustrated in the first row. Experts in the lower 16 layers tend to be more active than those in the upper layers. Among the experts, LoRA(Q, K) and Prompt experts show higher activation than LoRA(O, V, UP) and the Parallel Adapter. This indicates that MoA is capable of leveraging the distinct adaptation capabilities of heterogeneous experts according to their structural characteristics and positions within the model. In contrast, as shown in the second row, AdaMoLE fails to exhibit such consistency under different seeds, aligning with findings from prior work such as HMoE (Wang et al., 2024), which identifies under-specialization as a common issue in homogeneous expert settings.
@@ 6.3 Expert Activation Comparison
6.4 Instance- and Token-level Routing
In contrast to token-level routing approaches, instance-level routing, wherein the model makes a single routing decision for the entire input example, significantly reduces the computational overhead associated with the router. This brings the overall computational cost closer to that of standard, non-MoE architectures. We conducted a comparison between token-level MoA and instance-level variant of Soft MoA and the UniPEFT approach. The details are shown in Appendix B.1.
@@ Appendix A PEFT Methods And The Unified Form
Parallel Adapters involve adding small adapter modules in parallel to the Feed-Forward Network (FFN) or Attention blocks within Transformer layers. The adapter module is normally moulded by a two-layer feed-forward neural network with a bottleneck: a down-projection with $\boldsymbol{W}_{down}\in\mathbb{R}^{d\times r}$ to project the input $\boldsymbol{h}$ to a lower-dimensional space specified by bottleneck dimension $r$, followed by a nonlinear activation function $f(\cdot)$, an up-projection with $\boldsymbol{W}_{up}\in\mathbb{R}^{r\times d}$ to project back to the input size. The parallel adapter can be defined as:
@@ Appendix B Additional Experiments and Analyses
B.1 Comparison between Instance-level and Token-level Routing Methods
We conducted comparative experiments on LLaMA-3.1 8B, evaluating an instance-level variant of our Soft MoA method and the UniPEFT approach. UniPEFT utilizes a separate gated network for each type of adapter to control their influence on the input sample. As presented in Table 5, the results indicate that both instance-level approaches yield performance inferior to token-level routing methods. Furthermore, their performance is also surpassed by employing a standard LoRA adaptation alone.
@@ B.1 Comparison between Instance-level and Token-level Routing Methods
B.2 Case Study of Router Weight

### v2

@@ Abstract
Jie Cao1, Tianwei Lin 1, Bo Yuan 1, Rolan Yan 2, Hongyang He, Wenqiao Zhangthanks: Corresponding author 1, Juncheng Li 1, Dongping Zhang , Siliang Tang 1, Yueting Zhuang 1 1 Zhejiang University, 2 Tencent {caojie, lintw, byuan, wenqiaozhang, junchengli, siliang, yzhuang}@zju.edu.cn, rolanyan@qq.com
@@ 1 Introduction
The rapid development of Large Language Models (LLMs) (Achiam et al., 2023; Yang et al., 2025; Grattafiori et al., 2024a) is reshaping the field of NLP, demonstrating cross-domain generalization capabilities across a wide range of tasks. However, as model sizes continue to grow, the computational, storage, and deployment costs of traditional full fine-tuning methods increasingly become bottlenecks in practical applications. As a result, parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; He et al., 2022; Li and Liang, 2021; Lester et al., 2021; Zhang et al., 2024; Han et al., 2024; Lin et al., 2025) are emerging as a key research direction for adapting LLMs. These methods introduce a small number of trainable, lightweight adapter modules on top of frozen pre-trained weights. By doing so, they effectively guide the activation and adjustment of the model’s existing knowledge representations while maintaining low overhead, enabling fast transfer and adaptation to specific downstream tasks.
@@ 3.2 Sparse MoA
As the core objective of PEFT methods is to fully leverage LLMs’ pre-trained knowledge and capabilities via adapters for downstream tasks (Ghosh et al., 2024), Sparse MoA enhances this efficiency by adaptively activating diverse combinations of heterogeneous experts. Moreover, Sparse MoA adjusts the number of participating experts based on the contextual importance of tokens, thereby mitigating redundant expert computations. In extreme cases where certain tokens require no expert intervention and can be processed solely by the frozen LLM, Sparse MoA deactivates all experts, thus maximizing computational efficiency.
@@ 4.1 Datasets
We conduct extensive experiments and analyses on mathematical and commonsense reasoning tasks. Details for code generation tasks are provided in Appendix B.4. For mathematical reasoning, we evaluate performance on six test datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MultiArith Roy and Roth (2016), AddSub Hosseini et al. (2014), AQuA Ling et al. (2017), and SingleEq Koncel-Kedziorski et al. (2015). The training dataset is Math14K constructed by Hu et al. (2023), which includes the training sets of GSM8K and AQuA, augmented with rationales generated by ChatGPT and GPT-4. For commonsense reasoning, we test on eight datasets: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag, WinoGrande Sakaguchi et al. (2021), ARC Challenge (ARC-c), ARC Easy (ARC-e) Chollet (2019), and OBQA Mihaylov et al. (2018). The training dataset for commonsense reasoning, Commonsense15K, is sampled from Commonsense170K Hu et al. (2023) which comprises eight training sets.
@@ 4.2 Baseline Models
To validate the efficacy of our method, we compare it against: (i) the standard LoRA(Hu et al., 2021); (ii) SoTA soft-weighted parameter-efficient MoE methods: MoLoRA Zadouri et al. (2023) and HydraLoRA Tian et al. (2024); (iii) SoTA sparse parameter-efficient MoE approaches: MoLA (Gao et al., 2024) and AdaMoLE Liu and Luo (2024); (iv) other LoRA-based methods DenseLora Mu et al. (2025) and FlyLoRA Zou et al. .
@@ 4.3 Experiment Settings
Our experiments use the LLaMA3.1-8B model (Grattafiori et al., 2024b), Qwen3-8B model, and the larger Qwen3-14B model Yang et al. (2025). To ensure fair comparisons, we apply identical configurations to all LoRA-based methods: LoRA is inserted into four weight matrices in the self-attention module ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$) and one in the FFN module ($W_{\text{down}}$). For our MoA methods and all MoE-LoRA baselines, we set the number of experts to 8, with both the rank and alpha fixed at 8. To ensure a fair comparison with our MoA methods, we adjust the ranks of LoRA, FlyLoRA, and DenseLoRA to 16, 32, and 256, respectively, to maintain an equivalent number of trainable parameters. MoA integrates 7 heterogeneous experts, including five LoRA modules, parallel adapters in the FFN layer, and a zero-initialized prompt-tuning module. The bottleneck dimension of the parallel adapter is set to 16, and the prompt length to 10. We use the AdamW optimizer with a learning rate of 6e-3. For commonsense reasoning, input sequences are truncated to 200 tokens; for mathematical reasoning, to 300 tokens. All models are trained on A6000-48G GPUs.
The experimental results on Qwen3-8B and the larger Qwen3-14B are given in the appendix B.3.
@@ 5.1 Ablation Study on Soft MoA Components
LoRA+Prompt+PA 81.58 78.28
LoRA+Prompt+PA Softmax R 84.16 80.21
Soft MoA Sigmoid R 84.96 81.51
@@
The superior performance of LoRA over Prompt-tuning and Parallel Adapter, combined with its inherent inclusion of positionally diverse expert modules, motivated our extension of LoRA’s multi-module experts with a token-level router. While this adaptation yielded significant improvements over standalone LoRA (e.g., +1.32 on mathematical reasoning), it underperformed the full MoA (-1.01), highlighting two critical insights: 1. MoA’s primary strength derives from leveraging multiple LoRA modules as heterogeneous experts. 2. The distinct mechanisms of Prompt Tuning and Parallel Adapter further complement these LoRA-based experts.
@@ 6 In-depth Analysis
Figure 3: Comparison of router weight distributions between Soft MoA and AdaMoLE on the BoolQ test set under different random seeds. The MoA method exhibits strong consistency, whereas AdaMoLE does not.
@@ 6.2 Difference between MoA and MoE
As shown in Figure 3, we compare the expert weight distributions across 32 Transformer layers on 50 test samples for Soft MoA and AdaMoLE, trained with different random seeds. Soft MoA exhibits highly consistent expert weight distributions across seeds, as illustrated in the first row. Experts in the lower 16 layers tend to be more active than those in the upper layers. Among the experts, LoRA(Q, K) and Prompt experts show higher activation than LoRA(O, V, Down) and the Parallel Adapter. This indicates that MoA is capable of leveraging the distinct adaptation capabilities of heterogeneous experts according to their structural characteristics and positions within the model. In contrast, as shown in the second row, AdaMoLE fails to exhibit such consistency under different seeds, aligning with findings from prior work such as HMoE (Wang et al., 2024), which identifies under-specialization as a common issue in homogeneous expert settings.
Inter-expert activation similarity.
To analyze the architectural differences between MoA and homogeneous MoE, we measure expert specialization using the inter-expert activation standard deviation ($\sigma_{act}$). This metric is computed per token across all layers and subsequently averaged. A higher $\sigma_{act}$ indicates distinct expert specialization, whereas a lower value signifies "representation collapse"—a state where experts exhibit high redundancy and respond nearly identically to inputs. The metric is defined as:
$\sigma_{act}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(a_{i}-\bar{a})^{2}},$
where $N$ is the number of experts, $a_{i}$ is the activation of expert $E_{i}$, and $\bar{a}$ is the mean activation across all experts.
  Model Soft MoA Sparse MoA AdaMoLE
$\sigma_{act}$ ($\uparrow$) 0.2330 0.2769 0.0643
Table 4: Comparison of Inter-expert Activation Standard Deviation between MoA and AdaMoLE on the AddSub Dataset.
The results in the table 4 highlight a significant divergence between MoA and AdaMoLE. The minimal inter-expert activation standard deviation in AdaMoLE suggests a representation collapse, where experts fail to specialize. Conversely, MoA maintains a high degree of specialization, a benefit we attribute to its heterogeneous expert architecture.
@@ Appendix A PEFT Methods And The Unified Form
Parallel Adapters involve adding small adapter modules in parallel to the Feed-Forward Network (FFN) or Attention blocks within Transformer layers. The adapter module is normally moulded by a two-layer feed-forward neural network with a bottleneck: a down-projection with $\boldsymbol{W}_{down}\in\mathbb{R}^{d\times r}$ to project the input $\boldsymbol{x}$ to a lower-dimensional space specified by bottleneck dimension $r$, followed by a nonlinear activation function $f(\cdot)$, an up-projection with $\boldsymbol{W}_{up}\in\mathbb{R}^{r\times d}$ to project back to the input size. The parallel adapter can be defined as:
@@ Appendix B Additional Experiments and Analyses
B.1 Effect of Threshold Function
  Max Gamma Threshold Function Math
0.2 81.22
0.5 80.92
0.8 78.58
0.2 81.24
0.5 81.51
0.8 81.36
1 80.61
B.2 Comparison between Instance-level and Token-level Routing Methods
In contrast to token-level routing approaches, instance-level routing, wherein the model makes a single routing decision for the entire input example, significantly reduces the computational overhead associated with the router. This brings the overall computational cost closer to that of standard, non-MoE architectures. We conducted a comparison between token-level MoA and the instance-level variant of Soft MoA and the UniPEFT approach on LLaMA3.1-8B. UniPEFT utilizes a separate gated network for each type of adapter to control their influence on the input sample. As presented in Table 6, the results indicate that both instance-level approaches yield performance inferior to token-level routing methods. Furthermore, their performance is also surpassed by employing a standard LoRA adaptation alone.
@@ B.2 Comparison between Instance-level and Token-level Routing Methods
B.3 Foundation Model Analysis
To validate the effectiveness of our method across different and larger foundation models, we conducted experiments on Qwen3-8B and the larger-scale Qwen3-14B Yang et al. (2025). For the experiments on Qwen3-8B, we adopted the same experimental settings as LLaMA-3.1 8B, as described in Section 4.3. However, when using Qwen3-14B, we encountered GPU Out-of-Memory (OOM) issues for MoLoRA, HydraLoRA, and AdaMoLE when the number of experts was set to 8. Consequently, for these three methods, we set the number of experts to 4. In contrast, for MoLA, we still maintained a total of 8 experts, and only activated the top 2.
As detailed in Tables 7 and  8, our proposed Soft MoA and Sparse MoA methods outperform both LoRA and other MoE-LoRA approaches on mathematical reasoning benchmarks. Specifically, on the Qwen3-8B model, Soft MoA surpasses competing methods that utilize 2 to 4 times more parameters. On the larger Qwen3-14B model, Sparse MoA achieves the highest performance among all baselines while using the fewest parameters (34.50M). These results collectively demonstrate the superior parameter efficiency of the heterogeneous MoA approach compared to conventional MoE-LoRA methods, and further affirm its robustness across different foundation models and scales.
  Models AddSub AQuA GSM8k MultiArith SingleEq SVAMP Average Param
LoRA $93.67_{\pm 0.39}$ $37.40_{\pm 2.58}$ $81.80_{\pm 0.67}$ $98.17_{\pm 0.86}$ $96.85_{\pm 0.23}$ $88.80_{\pm 0.53}$ $82.78_{\pm 0.47}$ 24.77M
MoLoRA $93.76_{\pm 1.30}$ $42.26_{\pm 2.41}$ $82.51_{\pm 2.35}$ $98.11_{\pm 0.38}$ $96.06_{\pm 1.04}$ $90.40_{\pm 0.30}$ $\underline{83.85}_{\pm 1.17}$ 107.35M
HydraLoRA $93.16_{\pm 1.27}$ $39.11_{\pm 2.79}$ $81.78_{\pm 1.48}$ $98.11_{\pm 0.25}$ $95.21_{\pm 0.75}$ $88.80_{\pm 0.26}$ $82.70_{\pm 0.88}$ 49.55M
MoLA $93.33_{\pm 0.89}$ $42.78_{\pm 0.45}$ $81.50_{\pm 1.14}$ $98.33_{\pm 0.44}$ $96.13_{\pm 0.50}$ $89.83_{\pm 1.16}$ $83.65_{\pm 0.26}$ 107.35M
AdaMoLE $92.24_{\pm 1.05}$ $41.73_{\pm 4.26}$ $81.65_{\pm 1.00}$ $98.22_{\pm 0.25}$ $96.19_{\pm 0.80}$ $89.10_{\pm 0.75}$ $83.19_{\pm 0.85}$ 108.38M
FlyLoRA $93.00_{\pm 0.81}$ $39.11_{\pm 1.98}$ $86.20_{\pm 0.50}$ $98.44_{\pm 0.10}$ $95.87_{\pm 0.86}$ $89.23_{\pm 0.70}$ $83.64_{\pm 0.40}$ 24.77M
Soft MoA $93.50_{\pm 0.15}$ $43.31_{\pm 4.09}$ $85.17_{\pm 1.03}$ $98.06_{\pm 0.51}$ $96.72_{\pm 0.50}$ $88.30_{\pm 0.72}$ $\boldsymbol{84.17}_{\pm 0.89}$ 26.99M
Sparse MoA $93.33_{\pm 0.15}$ $40.68_{\pm 2.95}$ $83.85_{\pm 0.35}$ $98.17_{\pm 0.44}$ $97.31_{\pm 0.11}$ $88.80_{\pm 1.66}$ $83.69_{\pm 0.76}$ 24.49M
Table 7: Experimental results on Qwen3-8B for mathematical reasoning benchmarks. The evaluation metric is accuracy. All methods are run with three random seeds; mean and standard deviation are reported. “Param” indicates trainable parameters.
  Models AddSub AQuA GSM8k MultiArith SingleEq SVAMP Average Param
LoRA $96.20_{\pm 0.25}$ $41.34_{\pm 3.39}$ $87.49_{\pm 0.86}$ $97.83_{\pm 0.29}$ $97.64_{\pm 0.41}$ $90.60_{\pm 0.56}$ $85.18_{\pm 0.86}$ 35.39M
MoLoRA $96.37_{\pm 0.29}$ $44.88_{\pm 2.73}$ $87.59_{\pm 1.12}$ $98.33_{\pm 0.29}$ $97.90_{\pm 0.45}$ $91.50_{\pm 0.69}$ $86.10_{\pm 0.72}$ 76.84M
HydraLoRA $95.78_{\pm 1.39}$ $43.70_{\pm 2.39}$ $88.05_{\pm 1.20}$ $98.17_{\pm 1.04}$ $97.77_{\pm 0.41}$ $91.37_{\pm 0.81}$ $85.81_{\pm 0.34}$ 40.47M
MoLA $96.96_{\pm 0.67}$ $42.26_{\pm 4.55}$ $88.98_{\pm 0.68}$ $98.67_{\pm 0.17}$ $97.83_{\pm 0.52}$ $91.63_{\pm 0.57}$ $86.06_{\pm 0.66}$ 153.68M
AdaMoLE $96.20_{\pm 0.25}$ $40.55_{\pm 4.54}$ $88.02_{\pm 1.26}$ $98.50_{\pm 0.17}$ $98.10_{\pm 0.23}$ $91.13_{\pm 0.70}$ $85.42_{\pm 0.54}$ 78.36M
FlyLoRA $95.27_{\pm 0.15}$ $41.86_{\pm 2.24}$ $88.53_{\pm 0.52}$ $98.11_{\pm 0.35}$ $96.72_{\pm 0.50}$ $91.47_{\pm 0.15}$ $85.33_{\pm 0.40}$ 35.39M
Soft MoA $96.96_{\pm 0.25}$ $44.88_{\pm 2.98}$ $88.78_{\pm 0.54}$ $98.33_{\pm 0.10}$ $98.62_{\pm 0.80}$ $90.80_{\pm 0.25}$ $\boldsymbol{86.40}_{\pm 0.72}$ 37.98M
Sparse MoA $96.29_{\pm 0.89}$ $44.62_{\pm 3.44}$ $88.20_{\pm 1.02}$ $98.11_{\pm 0.35}$ $98.10_{\pm 0.11}$ $91.63_{\pm 0.99}$ $\underline{86.16}_{\pm 0.77}$ 34.50M
Table 8: Experimental results on Qwen3-14B for mathematical reasoning benchmarks. The evaluation metric is accuracy. All methods are run with three random seeds; mean and standard deviation are reported. “Param” indicates trainable parameters.
B.4 Code generation tasks
Datasets and Evaluation
To evaluate the code generation performance of our method, we utilize the CodeAlpaca-20k Chaudhary (2023) dataset as our training set. The evaluation is conducted on the HumanEval Chen (2021) benchmark. We use the pass@$k$ metrics (where $k\in\{1,5,10\}$) to measure the performance on code generation tasks. The pass@$k$ denotes the probability of generating at least one functionally correct solution within $k$ attempts.
Experimental setting
For the code generation task, we employ Qwen3-8B and Llama3.1-8B as base models. We adopt hyperparameter configurations similar to those used for mathematical reasoning. Given that different baselines exhibit sensitivity to the number of training steps on the code generation task, we perform checkpoint selection using the validation set during training. Specifically, we train for 5 epochs on Qwen3-8B and 1 epoch on Llama3.1-8B. For code generation, the maxmium sequence length is set to 500.
Main results
As illustrated in Table 9, our method outperforms both MoE-LoRA and LoRA-based baselines (e.g., DenseLoRA and FlyLoRA) while utilizing the fewest parameters. Specifically, on Qwen3-8B, Sparse MoA surpasses LoRA by 10.61 points in pass@1 and exceeds the strongest baseline, HydraLoRA, by 3.41 points. We observe that sparse MoE-LoRA methods, such as MoLA and AdaMoLE, exhibit suboptimal performance on code generation tasks compared to soft-weighted variants like MoLoRA and HydraLoRA. Furthermore, while the performance of FlyLoRA fluctuates significantly between Llama 3.1-8B and Qwen3-8B, our MoA method maintains consistent superiority across both base models, demonstrating its robust generalizability.
  Llama3.1-8B Qwen3-8B
Methods Pass@1 Pass@5 Pass@10 # Parameters Pass@1 Pass@5 Pass@10 # Parameters
LoRA 26.22 54.41 67.07 23.06M 39.69 72.29 81.71 24.77M
MoLoRA 34.15 63.15 72.56 100.14M 43.78 73.49 82.93 107.35M
HydraLora 33.35 60.96 69.51 45.09M 46.89 75.61 84.15 49.55M
MoLA 28.72 56.53 67.68 100.14M 41.10 71.60 81.10 107.35M
AdaMoLE 21.52 44.58 56.10 101.12M 40.91 66.65 76.83 108.38M
DenseLoRA 31.58 58.71 67.07 26.74M 46.82 76.04 84.75 27.00M
FlyLoRA 32.32 63.21 71.95 23.06M 20.18 38.15 48.17 24.77M
SoftMoA 31.10 59.86 70.73 24.52M 50.61 80.57 89.02 26.70M
SparseMoA 35.55 63.26 70.73 22.29M 50.30 81.12 87.81 24.49M
Table 9: Experimental results for the code generation task on Llama3.1-8B and Qwen3-8B. The evaluation metric is pass@k. "# Parameters" indicates trainable parameters. Bold numbers indicate the highest performance scores and underline numbers indicate the second performance scores.
B.5 Case Study of Router Weight